# An interpretable approach to automating the assessment of biofouling in video footage

### v1

@@ 2.4 Application to video footage
A stream of still images can be extracted from video footage, allowing a model only trained to assess single frames to be used to identify key features in a video. In this work, the challenge of identifying the SLoF level present in an image is defined on the basis of a single frame, so it is natural for videos to also be assessed in this way. However, this introduces additional complexities, as videos can be made up of a significant number of frames, and neural networks have been found to be sensitive to the perturbations between neighbouring frames within a video stream and predictions can be noisy.[22] To address this, the models are evaluated on only every $n$th frame to reduce computational costs, and Gaussian kernel smoothing[23] is used to average the predicted confidence and coverage of biofouling over neighbouring frames. This latter approach has the effect of stabilising predictions over time, and an example of smoothed versus unsmoothed outputs is shown in appendix D in the supporting information.
@@
Footage may also contain a large amount of imagery that is not relevant to identifying the presence of biofouling, such as video of open water and on-dock activities. To filter out these non-relevant sections, a further hull detection model is trained using ComFe as described in appendix C in the supporting information. A version of these models suitable for deployment is available for download from figshare, as highlighted on GitHub.[20]
@@ 3.1 ComFe models are interpretable
In fig. 1 an example ComFe explanation is shown, which identifies the image regions which result in a fouling present classification being made, and highlights exemplars from the training dataset that support this conclusion. The full set of class prototypes and their exemplars is shown in appendix A in the supporting information. Further examples are shown in fig. 3, demonstrating the ability of the ComFe model to identify where in an image biofouling occurs when making a prediction, without being explicitly trained with segmentation annotations.
@@ Acknowledgements
We would like to thank everyone that provided data towards this project. We would like to thank Ian Davidson, Oliver Floerl, Shaun Cunningham and Katherine Walls from Cawthron and New Zealand’s Ministry of Primary Industries. We would like to thank Ashley Coutts from Biofouling Solutions. We would like to thank Sarah Bailey from Fisheries and Oceans Canada. We would like to thank Lina Ceballos and Chris Scianni from the California State Lands Commision. We would like to thank Kevin Ellard from the Tasmanian Department of Primary Industries. We would like to thank Abagail Robinson, Daniel Kluza and Charlie Parker from the New Zealand’s Ministry of Primary Industries.
@@
Data and model weights are available from figshare.[20] Code for training ComFe models is available on GitHub.[15]
@@
B.W. collected the ROV video data, and led the gathering of data from other organisations. E.M. developed the methodology, applied it to the data, analysed the results and wrote the manuscript. All authors reviewed the manuscript.

### v2

@@ 2.4 Application to video footage
A stream of still images can be extracted from video footage, allowing a model only trained to assess single frames to be used to identify key features in a video. In this work, the challenge of identifying the SLoF level present in an image is defined on the basis of a single frame, so it is natural for videos to also be assessed in this way. However, this introduces additional complexities, as videos can be made up of a significant number of frames, and neural networks have been found to be sensitive to the perturbations between neighbouring frames within a video stream and predictions can be noisy.[22] To address this, the models are evaluated on only every $n$th frame to reduce computational costs, and Gaussian kernel smoothing[23] is used to average the predicted confidence and coverage of biofouling over neighbouring frames. This latter approach has the effect of stabilising predictions over time, and an example of smoothed versus unsmoothed outputs is shown in the supporting information.
@@
Footage may also contain a large amount of imagery that is not relevant to identifying the presence of biofouling, such as video of open water and on-dock activities. To filter out these non-relevant sections, a further hull detection model is trained using ComFe as described in the supporting information. A version of these models suitable for deployment is available for download from figshare, as highlighted on GitHub.[20]
@@ 3.1 ComFe models are interpretable
In Fig. 1 an example ComFe explanation is shown, which identifies the image regions which result in a fouling present classification being made, and highlights exemplars from the training dataset that support this conclusion. The full set of class prototypes and their exemplars is shown in the supporting information. Further examples are shown in Fig. 3, demonstrating the ability of the ComFe model to identify where in an image biofouling occurs when making a prediction, without being explicitly trained with segmentation annotations.
@@ Acknowledgements
We would like to thank everyone that provided data towards this project. We would like to thank Ian Davidson, Oliver Floerl and Shaun Cunningham from Cawthron, and Samantha Happy from Auckland Council. We would like to thank Ashley Coutts from Biofouling Solutions. We would like to thank Sarah Bailey from Fisheries and Oceans Canada. We would like to thank Lina Ceballos and Chris Scianni from the California State Lands Commision. We would like to thank Kevin Ellard from the Tasmanian Department of Primary Industries. We would like to thank Abagail Robinson, Daniel Kluza, Charlie Parker and Katherine Walls from the New Zealand’s Ministry of Primary Industries.
@@
Data and model weights are available from doi.org/10.6084/m9.figshare.26537158.v4. Code for training ComFe models is available at github.com/emannix/comfe-component-features.
@@
E.M. developed the methodology, applied it to the data, analysed the results and wrote the manuscript. B.W. collected the ROV video data, advised on project direction and led the gathering of data from other organisations. All authors reviewed the manuscript.